# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v1

@@ 5.3. Future Research Directions
On the security front, addressing various types of attacks in FL-based SG systems is paramount. Byzantine attacks, for instance, pose a serious threat as malicious clients can send incorrect model updates, corrupting the global model. These attacks are particularly challenging because they blend in with honest client contributions, making them difficult to detect. Techniques such as robust aggregation methods (e.g., Krum and Bulyan) have been proposed, but further refinement is needed to protect FL systems from Byzantine failures. Backdoor attacks are another prominent threat in which adversaries implant hidden malicious behaviors into the global model while still achieving high performance on regular tasks. This is especially concerning in SG applications, where unnoticed tampering could lead to severe disruptions. Enhancing defense mechanisms such as differential privacy and anomaly detection algorithms can mitigate the risk of backdoor attacks. Inference attacks, where adversaries attempt to extract sensitive information from the shared model, are also a critical area for research. While FL inherently provides some protection by not sharing raw data, adversaries can still exploit the model’s updates to infer private details. Improving privacy-preserving techniques, such as homomorphic encryption and secure multi-party computation, can help protect against these inference attacks. To defend FL-based SG systems from these various threats, several advanced defense mechanisms are being developed. Byzantine-resilient aggregation methods like Trimmed-Mean (Yin et al., 2018b), Krum (Blanchard et al., 2017), and Bulyan (Guerraoui et al., 2018) offer robust solutions by filtering out malicious model updates. These approaches, however, need further exploration in the context of SGs, where the scale and complexity of the system introduce unique challenges. FedVRDP (Zhang and Hu, 2023) is another promising approach that combines differential privacy with Byzantine resilience, offering a dual-layered defense against both privacy breaches and malicious attacks. Research into these and other defense strategies will be essential to making FL a viable and secure solution for real-world SG applications.

### v2

@@ 5.3. Future Research Directions
On the security front, addressing various types of attacks in FL-based SG systems is paramount. Byzantine attacks, for instance, pose a serious threat as malicious clients can send incorrect model updates, corrupting the global model. These attacks are particularly challenging because they blend in with honest client contributions, making them difficult to detect. Techniques such as robust aggregation methods (e.g., Krum and Bulyan) have been proposed, but further refinement is needed to protect FL systems from Byzantine failures. Backdoor attacks are another prominent threat in which adversaries implant hidden malicious behaviors into the global model while still achieving high performance on regular tasks. This is especially concerning in SG applications, where unnoticed tampering could lead to severe disruptions. Enhancing defense mechanisms such as differential privacy and anomaly detection algorithms can mitigate the risk of backdoor attacks. Inference attacks, where adversaries attempt to extract sensitive information from the shared model, are also a critical area for research. While FL inherently provides some protection by not sharing raw data, adversaries can still exploit the model’s updates to infer private details. Improving privacy-preserving techniques, such as homomorphic encryption and secure multi-party computation, can help protect against these inference attacks. To defend FL-based SG systems from these various threats, several advanced defense mechanisms are being developed. Byzantine-resilient aggregation methods like Trimmed-Mean (trmean_median), Krum (krum), and Bulyan (bulyan) offer robust solutions by filtering out malicious model updates. These approaches, however, need further exploration in the context of SGs, where the scale and complexity of the system introduce unique challenges. FedVRDP (zhang2023byzantine) is another promising approach that combines differential privacy with Byzantine resilience, offering a dual-layered defense against both privacy breaches and malicious attacks. Research into these and other defense strategies will be essential to making FL a viable and secure solution for real-world SG applications.